A triglyceride contains lauric acid (12:0), linoleic acid (18:2) and palmitoleic acid (16:1) chains. How many moles of H2 would be required to completely reduce this triglyceride to a saturated molecule?
Answer #1

Lauric acid is saturated fatty acid so that does not need any hydrogen for saturation

Lauric acid (12:0) :    CH3-CH2-(-CH2)9-COOH 12 represents number of carbon atoms and 0 represents number of double bonds

Linoleic acid is unsaturated fatty acid. Linoleic acid (18:2) : 18 represents number of carbon atoms and 2 represents number of double bonds (unsaturation)

  CH3-(CH2)4-CH=CH-CH2-CH=CH-(CH2)7-COOH

Palmitoleic acid (16:1) : It is also unsaturated fatty acid 16 represents number of carbon atoms and 1 represents number of double bonds (unsaturation).

CH3-(CH2)5 -CH=CH-(CH2)7-COOH

For one double bond one mole of hydrogen (H2) is required. As the triglyceride contains auric acid (12:0), linoleic acid (18:2) and palmitoleic acid (16:1) chains.

Totally there are 3 unsaturated bonds. Hence 3 moles of hydrogen (H2) are required to completely reduce this triglyceride to a saturated molecule.